mikewwe13:

What is the average rate of change of the function f (x) = 4(2)^x from x = 0 to x = 2? Enter your answer in the box. ____________

mikewwe13:

@Vocaloid

Vocaloid:

if you didn't understand something about moreso's work, you are welcome to ask questions

mikewwe13:

i didn't know how to do it and it's better if you teach me

Vocaloid:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @moreso \(\color{#0cbb34}{\text{Originally Posted by}}\) @moreso \(\color{#0cbb34}{\text{Originally Posted by}}\) @moreso use the formula \[ \Large f_{avg}= \frac{f(b)-f(a)}{b-a}\] \(\color{#0cbb34}{\text{End of Quote}}\) \(\color{#0cbb34}{\text{End of Quote}}\) \(\color{#0cbb34}{\text{End of Quote}}\) in this case, b is your "upper" x value (2) and a is your lower x value (0) so let b = 2 and a = 0 and calculate the rate using the above formula

mikewwe13:

i just calculated it and its 40 ?

moreso:

thats incorrect

moreso:

\[ \large f_{avg}= \frac{ 4(2)^2-4(2)^0}{2-0}\]

Vocaloid:

i kind of need to get work done tonight, I'll let moreso take over ^_^ thanks

moreso:

\[ \large f_{avg}= \frac{ 4(2)^2-4(2)^0}{2-0} = \frac{4\times 4 - 4 \times 1 }{2} \]

mikewwe13:

14

moreso:

(16-4)/2

mikewwe13:

6

mikewwe13:

it's 6

mikewwe13:

correct me if i'm wrong

moreso:

yes 6

mikewwe13:

ok thanks moreso and vocaloid